OpenStudy (anonymous):

log[4](xy)^3 - log[4](xy) Simplify this

OpenStudy (apoorvk):

\[\log \left(\begin{matrix}m \\ n\end{matrix}\right) - \log \left(\begin{matrix}n \\a\end{matrix}\right) = \log \left(\begin{matrix}m \div n \\ a\end{matrix}\right)\]

OpenStudy (king):

can u give the equation using equation editor...

OpenStudy (anonymous):

\[\LARGE \log_{4}(xy)^3-\log_{4}(xy)=\log_{4}\left(\frac{x^3y^3}{xy}\right)\]

OpenStudy (anonymous):

continue...

OpenStudy (anonymous):

\[\log((xy)^4)-\log(xy)=\log(\frac{x^4y^4}{xy})=\log(x^3y^3)=\log((xy)^3)=3\log(xy)\] is one method

OpenStudy (anonymous):

simpler method is \[\log((xy)^4)-\log(xy)=4\log(xy)-\log(xy)=3\log(xy)\]

OpenStudy (anonymous):

silly me it is \[\log((xy)^3)\]so answer is \[2\log(xy)\]

OpenStudy (anonymous):

log[4](xy)^3 - log[4](xy) =3log[4]x + 3log[4]y - (log[4]x + log[4] y) =2log[4]x + 2log[4]y =2log[4](xy) =log[4](xy)^2 is this correct?